OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point. y = sec x, (π/6, 2 3 /3)

hartnn (hartnn):

y= sec x can you first find dy/dx ?

OpenStudy (anonymous):

y=secx tanx

hartnn (hartnn):

you mean dy/dx = sec x tan x, right ? now just plug in values of your point x= pi/6

OpenStudy (anonymous):

I got 2x+2rt3/3-pi/3 but it is incorrect apparently

hartnn (hartnn):

dy/dx = slope = sec (pi/6) tan (pi/6) = 2/3 did you get this first ?

OpenStudy (anonymous):

Shouldn't it be 2?

hartnn (hartnn):

sec pi/6 = 2/sqrt 3 tan pi/6 = 1/sqrt 3 so, its 2/ sqrt3 * (1/sqrt 3) = 2/3

hartnn (hartnn):

still confused ?

OpenStudy (anonymous):

Yes because I tried that and it still did not work

hartnn (hartnn):

thats just slope, we still have to find the equation of line...

hartnn (hartnn):

we have slope= m= 2/3 now and a point, (pi/6, 2 3/3) = (x1,y1) use the slope point form \(y- y_1 = m (x-x_1)\)

OpenStudy (anonymous):

I know and I plugged that in and tried that answer which did not work

hartnn (hartnn):

whats your y-co-ordinate in the question ?

OpenStudy (anonymous):

2rt3/3

hartnn (hartnn):

ok, so its \(2\sqrt 3/3\) y- \(2\sqrt 3/3 = (2/3) (x -\pi/6)\) \(3y -2\sqrt 3 = 2x - \pi/3\)

hartnn (hartnn):

so, \(3y -2x = 2\sqrt3 -\pi/3\) try this.

OpenStudy (anonymous):

It has to be y=mx+b

hartnn (hartnn):

okk, \(3y = 2x +2\sqrt 3 -\pi/3 \\ y= (2/3 )x + (2/\sqrt 3 -\pi/9)\)

OpenStudy (anonymous):

Just got it. Thank you.

hartnn (hartnn):

welcome ^_^